# The Geometry of Grokking: Norm Minimization on the Zero-Loss Manifold

### v1

@@ 1 Introduction
Neural networks have achieved great success, but their mechanisms remain far from being fully understood. Doshi-Velez & Kim (2017) argue that understanding the inner workings of neural networks is crucial for the development of AI systems with increased safety and reliability. Moreover, understanding the learning dynamics of neural networks could also help us improve their performance and efficiency: it is easier to design better learning algorithms when we understand the limitations of the existing ones. Furthermore, insights into artificial neural networks may also enhance our understanding of biological neural networks due to their fundamental similarities (Sucholutsky et al., 2023; Kohoutová et al., 2020).
@@
Moreover, the role of the embedding layer in the modular addition task is a striking example that that generalization in neural networks often hinges on representation learning within specific network components. In such cases, it can be highly beneficial to simplify the complex learning dynamics of a deep network by isolating and analyzing only the component of interest. Recent work has already begun to explore such approximations, also referred to as effective theories (van Rossem & Saxe, 2024; Liu et al., 2022a; Musat, 2024; Mehta et al., ).
@@ 3.2 Norm Minimization
While the previous example illustrates our theoretical framework, it does not capture its full generality. It is unsurprising that applying weight decay encourages a reduction in norm. However, the central claim of this paper is significantly stronger: we argue that the learning dynamics under weight decay do not merely follow some norm-decreasing direction, but rather evolve along the direction that optimally minimizes the norm, subject to remaining on the zero-loss manifold. Another way to view this is the following: once the model achieves perfect memorization, learning effectively follows gradient descent on the weight norm, constrained to the zero-loss manifold.
@@ 3.3 A Few Mathematical Nuances
An important caveat is that the zero-loss set is not necessarily a manifold everywhere: it might contain singular points. Such a singular point is demonstrated in Figure 3 (center). However, such singularities should not worry us too much. If the network is realized by a smooth function, then the singularities will form a null set, i.e. the probability of encountering a singularity during standard training is exactly zero.
@@ 4.1 Architecture
We consider a neural network trained with mean-squared error loss on $k$ samples:
$\mathcal{L}(\theta)=\sum_{i=1}^{k}\|f(\theta,x_{i})-y_{i}\|^{2}.$ (1)
where $x_{i}\in\mathbb{R}^{n}$, $y_{i}\in\mathbb{R}^{m}$, and $\theta\in\mathbb{R}^{d}$ is the parameter vector. The network has $d$ parameters, $n$ inputs, and $m$ outputs. We use $f:\mathbb{R}^{d\times n}\rightarrow\mathbb{R}^{m}$ to denote the network realization function.
We apply a weight decay term (Krogh & Hertz, 1991) with a coefficient $\lambda\in\mathbb{R}^{+}$ to obtain:
$\mathcal{L}_{\lambda}(\theta)=\mathcal{L}(\theta)+\lambda\|\theta\|^{2}.$ (2)
We also denote the combined outputs of all samples as using $\mathcal{F}(\theta)\in\mathbb{R}^{km}$, where
@@ 4.2 Theoretical Setup
The following concepts are fundamental to our analysis:
Definition 4.1 (Zero-Loss Set).
Definition 4.2 (Singular Points).
We say that $\theta\in\mathbb{R}^{d}$ is a singular point [^1] We define singular points in terms of $\mathcal{F}$, though they correspond to singular points of $\mathcal{Z}$. With respect to $\mathcal{F}$, they are more precisely critical points. if the Jacobian matrix of $\mathcal{F}$ at $\theta$ is not full rank, i.e. $\mathrm{rank}(\nabla\mathcal{F}(\theta))<\min(d,km)$.
@@
Assumption 4.5 (Gradient Flow).
We assume that $\mathcal{F}$ is smooth and we model the gradient descent trajectory as a continuous-time gradient flow:
$\frac{\partial\theta(t)}{\partial t}=-\nabla\mathcal{L}_{\lambda}(\theta(t))$ (4)
where $\theta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{d}$ is the parameter vector as a function of time.
@@
We study the learning dynamics in the approximation of a very small weight decay coefficient, i.e. $\lambda\rightarrow 0$, motivated by the small values of $\lambda$ typically used by practitioners (Smith, 2018).
Assumption 4.7 (Well-Behaved Loss).
We assume that our training trajectory does not pass through singular points. This is motivated by the fact that singular points form a null set, i.e. a set of Lebesgue measure zero (Morse, 1939; Sard, 1942).
4.3 Constrained to the Zero-Loss
@@ 4.4 Loss Gradient Orthogonality
We will now provide our main theoretical result, which states that $\nabla\mathcal{L}(\theta)$ around $\mathcal{Z}$ becomes orthogonal to any available direction.
Definition 4.10 (Available Direction).
We say that $v\in\mathbb{R}^{d}$ is an available direction at $\theta\in\mathcal{Z}$ if there exists a smooth trajectory $s:\mathbb{R}\rightarrow\mathbb{R}^{d}$ such that $s(0)=\theta$, $s^{\prime}(0)=v$, and $\mathcal{L}(s(t))=0$ for all $t\in\mathbb{R}$.
@@
We denote by $\mathcal{T}_{\theta}$ the set of all available directions at $\theta\in\mathcal{Z}$.
@@
Let $S\subseteq\mathbb{R}^{d}$ be a compact space with $\mathrm{proj}_{\mathcal{Z}}(S\,\cap\,\mathcal{Z})\subseteq S$. Assume that $S$ contains no singular points. Then, there exist constants $C>0$ and $x_{0}>0$ such that
$\left|\frac{\;\;v^{\top}}{\|v\|}\,\frac{\nabla\mathcal{L}(\theta)}{\|\nabla\mathcal{L}(\theta)\|}\right|\;<\;C\,\mathrm{dist}_{\mathcal{Z}}(\theta)$ (6)
holds for all $\theta\in S\setminus\mathcal{Z}$ with $\mathrm{dist}_{\mathcal{Z}}(\theta)<x_{0}$ and $v\in\mathcal{T}_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$.
@@
Using $v\in\mathcal{T}_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$ and the absence of singular points, we are able to show that $Hv=0$ and $\|\nabla\mathcal{L}(\theta)\|=\Theta(\|x\|)$. Therefore, the normalized dot product will be $O(\|x\|)$. We give the full proof in Appendix B. ∎
@@
In other words, the loss gradient does not induce any movement near $\mathcal{Z}$. After a memorizing solution is reached, learning will be driven entirely by weight decay. The loss will only serve to keep the model near $\mathcal{Z}$, while weight decay will be free to push the model towards norm minimization along any of the available directions.
@@ 5 Isolated Dynamics of a Network Component
where $\phi:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ is a function specific to the loss function and the initial parameters. This parametrization allows us to isolate the dynamics of $\theta_{1}$ by expressing them as a function of $\theta_{1}$ alone:
@@ 5.1 Approximate Cost Function
The learning dynamics of $\theta_{1}$ under Equations 9 and 10 follow the gradient flow of $\mathcal{R}$ when $\phi$ is differentiable:
@@ 6.2 Isolated Dynamics of the First Layer
Since the second layer is just a linear transformation of the hidden layer activations $H=\sigma(XW_{1})$, finding the optimal second layer weights is equivalent to the classic problem of ridge regression (Hoerl & Kennard, 1970). The solution is given by:
@@
By combining Equations 13, 14 and 16, we can obtain the cost function for the isolated learning dynamics of the first layer:
@@ 6.3 Zero-Loss Approximation
Following the theoretical framework developed in Section 4, we can further simplify equation Equation 16 by working in the limit of very small weight decay $\lambda\rightarrow 0$:
$\phi(W_{1})=(H^{\top}H)^{-1}H^{\top}Y$ (19)
The expression $H^{+}=(H^{\top}H)^{-1}H^{\top}$ is also known as the Moore-Penrose pseudo-inverse of $H$. If $H$ has full column rank, then the pseudo-inverse is also given by $H^{+}=H^{\top}(HH^{\top})^{-1}$. This allows us to further simplify the cost function down to:
@@
$\dot{W_{1}}\approx X^{T}\Bigl(\bigl(AYY^{T}AH\bigr)\odot\sigma^{\prime}\bigl(XW_{1}\bigr)\Bigr)-W_{1}$ (21)
where $H=\sigma(XW_{1})$, $A=(HH^{T})^{-1}$, $\sigma$ is the activation function, and $\odot$ denotes the Hadamard product. We provide a detailed derivation in Appendix C.
@@ 7.3 Simulated Optimization
We simulate the evolution of the embedding matrix $E$ under the isolated dynamics given by equation Equation 21. We start from a random initialization $E\sim\mathcal{N}(0,\,p^{-1/2})$ and update it for $T$ steps as $E\leftarrow E+\eta\,\Delta E$, where $\eta>0$ is the step size and $\Delta E=X^{T}\left(\left(AYY^{T}AH\right)\odot\sigma^{\prime}\bigl(XE\bigr)\right)-E$.
The isolated dynamics assume that $W=(H^{T}H)^{-1}H^{T}Y$ is optimal for the current value of $E$, which guarantees zero loss and perfect accuracy on the training data throughout training. This also ensures that predicted outputs perfectly match the target outputs on the training data, which is less principled for a classification task, but generally performs well in practice (Rifkin et al., 2003).
7.4 Details
@@ 8 Conclusion
We provide the full code for training and plotting used for the experiments with simulated dynamics from Section 7.
@@
Appendix B Proof of Theorem 4.13 (Gradient Orthogonality)
@@
Lemma B.1.
At a non-singular point of the zero-loss set, the tangent space is exactly the null space of the Hessian matrix, i.e. $\mathcal{T}_{\theta}=\{\;v\in\mathbb{R}^{d}\;|\;\nabla^{2}\mathcal{L}(\theta)\,v=0\;\}\$for any non-singular $\theta\in\mathcal{Z}$.
@@
Part I: $\nabla^{2}\mathcal{L}(\theta)\;v=0\;\Rightarrow\,v\in\mathcal{T}_{\theta}$.
@@
Theorem B.2 (Inverse Function Theorem).
Assume that $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{km}$ is a smooth function with $d\geq km$ and $\mathrm{rank}(\nabla\mathcal{F}(\theta))=km$ for some $\theta\in\mathbb{R}^{d}$. Then, for all $v\in R^{d}$ such that $\nabla\mathcal{F}(\theta)\,v=0$, there exists a smooth trajectory $s:R\rightarrow R^{d}$ such that $s(0)=\theta,\ s^{\prime}(0)=v,\,$ and $\mathcal{F}(s(t))=\mathcal{F}(\theta)\,$ for all $t\in\mathbb{R}$.
@@
Part II: $v\in\mathcal{T}_{\theta}\;\Rightarrow\,\nabla^{2}\mathcal{L}(\theta)\;v=0$.
@@
Definition B.3 (Normal Space).
Let $\mathcal{N}_{\phi}=\{\,\alpha\in\mathbb{R}^{d}\;|\;\alpha^{\top}v=0,\;\forall v\in\mathcal{T}_{\phi}\,\}$ denote the normal space at a point $\phi\in\mathcal{Z}$.
Proposition B.4.
@@
We give a proof by contradiction. Assume that there exists $v\in\mathcal{T}_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$ such that $v^{\top}\left(\theta-\mathrm{proj}_{\mathcal{Z}}(\theta)\right)\neq 0$. Then, by Definition 4.10, there must exist a smooth trajectory $f:\mathbb{R}\rightarrow\mathbb{R}^{d}$ with $f(0)=\mathrm{proj}_{\mathcal{Z}}(\theta),\ f^{\prime}(0)=v,$ and $f(t)\in\mathcal{Z}$ for all $t\in R$. By moving $\mathrm{proj}_{\mathcal{Z}}(\theta)$ along this trajectory, we can get closer to $\theta$. However, this should not be possible according to Definition 4.12. ∎
We now establish our main result: See 4.13
@@
with $\phi\in\mathcal{Z},\,\alpha\in\mathcal{N}_{\phi},\,x\in R^{+},$ and $v\in\mathcal{T}_{\phi}$.
@@
for all $\phi\in S\cap\mathcal{Z},\,\alpha\in U(\mathcal{N}_{\phi}),\,x>0,\,v\in\mathcal{T}_{\phi}\$, where $U(\mathcal{N}_{\phi})=\{\,v\in\,\mathcal{N}_{\phi}\;|\;\|v\|=1\,\}$.
@@
Since $v\in\mathcal{T}_{\phi}$, from Lemma B.1, we have that $v^{\top}\,\nabla^{2}\mathcal{L}(\phi)\,\alpha=0$. This gives
@@
for any $x<a_{\inf}$. An appropriate choice of $C$ and $x_{0}$ gives the desired bound, for example:

### v2

@@ 1 Introduction
Neural networks have achieved great success, but their mechanisms remain far from being fully understood. Doshi-Velez and Kim (2017) argue that understanding the inner workings of neural networks is crucial for the development of AI systems with increased safety and reliability. Moreover, understanding the learning dynamics of neural networks could also help us improve their performance and efficiency: it is easier to design better learning algorithms when we understand the limitations of the existing ones. Furthermore, insights into artificial neural networks may also enhance our understanding of biological neural networks due to their fundamental similarities (Sucholutsky et al., 2023; Kohoutová et al., 2020).
@@
Moreover, the role of the embedding layer in the modular addition task is a striking example that generalization in neural networks often hinges on representation learning within specific network components. In such cases, it can be highly beneficial to simplify the complex learning dynamics of a deep network by isolating and analyzing only the component of interest. Recent work has already begun to explore such approximations, also referred to as effective theories (van Rossem and Saxe, 2024; Liu et al., 2022a; Musat, 2024; Mehta et al., ).
@@ 3.2 Norm Minimization
While the previous example illustrates our theoretical framework, it does not capture its full generality. It is unsurprising that applying weight decay encourages a reduction in norm. However, the central claim of this paper is significantly stronger: we argue that the learning dynamics under weight decay do not merely follow some norm-decreasing direction, but rather evolve along the direction that maximally decreases the norm, subject to remaining on the zero-loss manifold. Another way to view this is the following: once the model achieves perfect memorization, learning effectively follows gradient descent on the weight norm, constrained to the zero-loss manifold.
@@ 3.3 A Few Mathematical Nuances
An important caveat is that the zero-loss set is not necessarily a manifold everywhere: it might contain singular points. Such a singular point is demonstrated in Figure 3 (center). However, such singularities should not worry us too much. As we prove in Theorem 4.10, if the network is realized by a smooth function, we will almost never encounter a singularity during standard training.
@@
3.4 Gradient Orthogonality
We further illustrate our key theoretical result using a toy loss landscape. Consider a model with only two parameters $x,y\in\mathbb{R}$ and a loss function $\mathcal{L}(x,y)=(y-x^{2})^{2}$. We visualize this loss landscape in Figure 4. We plot three points $A,B,C\in\mathbb{R}^{2}$ that have the same projection on the zero-loss manifold, but are getting progressively closer. As we prove in Theorem 4.14, the loss gradients become perfectly orthogonal to the zero-loss set as we approach it.
Figure 4: We illustrate Theorem 4.14 using a toy loss landscape $\mathcal{L}(x,y)=(y-x^{2})^{2}$. We plot the level sets of $\mathcal{L}$ from three different views. Each view shows a magnified region of the previous view, zooming into the zero-loss manifold. We display the gradient angle at three different points, pointed in the negative direction. The gradient norm is not shown to scale. Gradients become increasingly orthogonal to the zero-loss manifold as we get closer.
@@ 4.1 Architecture
We consider a neural network trained with mean-squared error loss on $k$ samples and weight decay:
$\mathcal{L}(\theta)=\sum_{i=1}^{k}\|f(\theta,x_{i})-y_{i}\|^{2}$ $\mathcal{L}_{\lambda}(\theta)=\mathcal{L}(\theta)+\lambda\|\theta\|^{2}$ (1)
where $x_{i}\in\mathbb{R}^{n}$, $y_{i}\in\mathbb{R}^{m}$, and $\theta\in\mathbb{R}^{d}$ is the parameter vector. The network has $d$ parameters, $n$ inputs, and $m$ outputs. We use $f:\mathbb{R}^{d\times n}\rightarrow\mathbb{R}^{m}$ to denote the network realization function. We apply a weight decay term (Krogh and Hertz, 1991) with a coefficient $\lambda>0$.
We also denote the concatenated outputs for all training samples using $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{km}$, where
@@ 4.2 Theoretical Setup
Assumption 4.2 (Smooth Network).
We assume that the network realization function $f$ is continuously differentiable dkm+1d-km+1 times.
Assumption 4.3 (Gradient Flow).
We model the gradient descent trajectory as a gradient flow. We consider the parameter vector as a continuous function of time $\theta:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{d}$ with dynamics:
$\frac{\partial\theta(t)}{\partial t}=-\nabla\mathcal{L}_{\lambda}(\theta(t)).$ (3)
@@
We study the learning dynamics in the approximation of a very small weight decay coefficient, i.e. $\lambda\rightarrow 0$, motivated by the small values of $\lambda$ typically used by practitioners (Smith, 2018), as well as by previous empirical works on grokking (Liu et al., 2022b) suggesting that a small $\lambda$ is essential for the delayed-generalization phenomona.
4.3 Constrained to the Zero-Loss Set
@@
4.4 Regularity of the Zero-Loss Set
We show that the zero-loss set is well-behaved for almost every dataset.
Definition 4.9 (Singular Points).
We say that $\theta\in\mathbb{R}^{d}$ is a singular point if the Jacobian matrix of $\mathcal{F}$ at $\theta$ is not full rank, i.e. $\mathrm{rank}(\nabla\mathcal{F}(\theta))<\min(d,km)$. We denote the set of all singular points as $\mathrm{Crit}(\mathcal{F})\subset\mathbb{R}^{km}$. Note that singular points are defined in terms of $\mathcal{F}$, though they correspond to singular points of $\mathcal{Z}$. With respect to $\mathcal{F}$, they are more precisely critical points.
Theorem 4.10 (Regularity of $\mathcal{Z}$).
For almost every dataset $(x_{i},y_{i})_{i\leq k}$, the corresponding zero-loss set $\mathcal{Z}$ will not contain any singular points.
Sketch of the Proof.
We show that for any set of input vectors $(x_{i})_{i\leq k}$, almost every possible set of target vectors $(y_{i})_{i\leq k}$ induces a zero-loss set $\mathcal{Z}$ that is completely free of singular points.
Using the assumption that $\mathcal{F}$ is smooth, our desired result follows almost immediately from Sard (1942), also known as the Morse–Sard theorem, which states that the image of the critical points has Lebesgue measure zero. In our case, the image of the singular points $\mathcal{F}\big(\mathrm{Crit}(\mathcal{F})\big)$ has Lebesgue measure zero in the space of possible outputs $\mathbb{R}^{km}$. In other words, only a negligible set of possible outputs is ever hit by a singular point. We give a detailed proof in Appendix B. ∎
@@ 4.5 Loss Gradient Orthogonality
We will now provide our main theoretical result, which states that $\nabla\mathcal{L}(\theta)$ around $\mathcal{Z}$ becomes orthogonal to any tangent direction. We illustrate this concept in Figure 4.
Definition 4.11 (Tangent Direction).
We say that $v\in\mathbb{R}^{d}$ is a tangent direction at $\theta\in\mathcal{Z}$ if there exists a smooth trajectory $s:\mathbb{R}\rightarrow\mathbb{R}^{d}$ such that $s(0)=\theta$, $s^{\prime}(0)=v$, and $\mathcal{L}(s(t))=0$ for all $t\in\mathbb{R}$.
@@
We denote by $T_{\theta}$ the set of all tangent directions at $\theta\in\mathcal{Z}$.
@@
Let $S\subset\mathbb{R}^{d}$ be a compact space with $\mathrm{proj}_{\mathcal{Z}}(S)\subseteq S$. If $S\cap\mathcal{Z}$ contains no singular points, then there exists a constant $C>0$ such that
$\bigg|\cos\Big(\angle\big(v,\nabla\mathcal{L}(\theta)\big)\Big.)\bigg|\;=\;\left|\frac{\;\;v^{\top}}{\|v\|}\,\frac{\nabla\mathcal{L}(\theta)}{\|\nabla\mathcal{L}(\theta)\|}\right|\;<\;C\,\mathrm{dist}_{\mathcal{Z}}(\theta)$ (5)
holds for all $\theta\in S\setminus\mathcal{Z}$ and all tangent directions $v\in T_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$.
@@
Using $v\in T_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$ and the absence of singular points, we are able to show that $Hv=0$ and $\|\nabla\mathcal{L}(\theta)\|=\Theta(\|x\|)$. Therefore, the normalized dot product will be $O(\|x\|)$. We give the full proof in Appendix C. ∎
@@
In other words, the loss gradient does not induce any movement near $\mathcal{Z}$. After a memorizing solution is reached, learning will be driven entirely by weight decay. The loss will only serve to keep the model near $\mathcal{Z}$, while weight decay will be free to push the model towards norm minimization along any of the tangent directions.
4.6 Empirical Validation
Figure 5: We train smalls networks to perform modular addition. We display their average train and test loss (left) and the average cosine similarity between the parameter updates and the norm-minimizing direction on the zero-loss set (right). As we can see from the loss plot (left), the network exhibits grokking. Interestingly, the similarity is greatest exactly during the grokking stage.
We empirically validate our theory that post-memorization dynamics follows the norm-minimization direction on the zero-loss manifold. We train a few small two-layer networks to perform modular addition with ReLU activation, mean squared-error loss, and weight decay. At every step $t$ during training, we measure the cosine similarity between the parameter update $\Delta\theta_{t}$ and an estimate of the norm-minimizing direction on the zero-loss set $\tilde{g}_{t}=\arg\min_{v}(v^{\top}\theta_{t})$ such that $v\in T_{\mathrm{proj}_{\mathcal{Z}}(\theta_{t})}$. We display the results in Figure 5. We explain the full experimental details in Appendix D.
@@ 5 Isolated Dynamics of a Network Component
where $\phi:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ is a function specific to the loss function and the initial parameters.
We motivate this assumption based on the comprehensive literature on multiple points of stochastic processes. For example, even with a dimensionality as small as $d\geq 4$, it is known that a Brownian motion in $\mathbb{R}^{d}$ contains no self-intersections almost surely (Mörters and Peres (2010), Chapter 9). More generally, Dalang et al. (2021) prove the non-existance of multiple points for a wide class of Gaussian random fields.
This parametrization allows us to isolate the dynamics of $\theta_{1}$ along the training trajectory by expressing them as a function of $\theta_{1}$ alone:
@@ 5.1 Approximate Cost Function
The learning dynamics of $\theta_{1}$ under Equations 8 and 9 follow the gradient flow of $\mathcal{R}$ when $\phi$ is differentiable:
@@ 6.2 Isolated Dynamics of the First Layer
Since the second layer is just a linear transformation of the hidden layer activations $H=\sigma(XW_{1})$, finding the optimal second layer weights is equivalent to the classic problem of ridge regression (Hoerl and Kennard, 1970). The solution is given by:
@@
By combining Equations 12, 13 and 15, we can obtain the cost function for the isolated learning dynamics of the first layer:
@@ 6.3 Zero-Loss Approximation
We further assume a strongly overparametrized regime with more hidden units that training samples ($d_{h}>n$). This allows the second layer to fit the outputs perfectly almost always. Following the theoretical framework developed in Section 4, we can further simplify equation Equation 15 by working in the limit of very small weight decay $\lambda\rightarrow 0$:
$\phi(W_{1})=H^{+}Y$ (18)
where $H^{+}$ is the Moore-Penrose pseudo-inverse of $H$. When $H$ has linearly independent columns, then the pseudo-inverse is given by $H^{+}=(H^{\top}H)^{-1}H^{\top}$. If $H$ has linearly independent rows, then $H^{+}=H^{\top}(HH^{\top})^{-1}$. Note that $H\in\mathbb{R}^{n\times d_{h}}$. Given the overparameterized regime ($d_{h}>n$), we use the latter. This allows us to further simplify the cost function down to:
@@
$\dot{W_{1}}\approx X^{\top}\Bigl(\bigl(AYY^{\top}AH\bigr)\odot\sigma^{\prime}\bigl(XW_{1}\bigr)\Bigr)-W_{1}$ (20)
where $H=\sigma(XW_{1})$, $A=(HH^{\top})^{-1}$, $\sigma$ is the activation function, and $\odot$ denotes the Hadamard product. We provide a detailed derivation in Appendix E.
@@ 7.3 Simulated Optimization
We simulate the evolution of the embedding matrix $E$ under the isolated dynamics given by equation Equation 20. We start from a random initialization $E\sim\mathcal{N}(0,\,p^{-1/2})$ and update it for $T$ steps as $E\leftarrow E+\eta\,\Delta E$, where $\eta>0$ is the step size and $\Delta E=X^{\top}\left(\left(AYY^{\top}AH\right)\odot\sigma^{\prime}\bigl(XE\bigr)\right)-E$.
The isolated dynamics assume that $W=(H^{\top}H)^{-1}H^{\top}Y$ is optimal for the current value of $E$, which guarantees zero loss and perfect accuracy on the training data throughout training. This also ensures that predicted outputs perfectly match the target outputs on the training data, which is less principled for a classification task, but generally performs well in practice (Rifkin et al., 2003).
Details.
@@ 8 Conclusion
We provide the full code for training and plotting used for the experiments from Sections 4.6, LABEL: and 7.
@@
Appendix B Proof of Theorem 4.10 (Regularity of $\mathcal{Z}$)
We consider the inputs vectors $x_{i}\in\mathbb{R}^{n}$ fixed and we show that singularities will not be encountered for almost all possible target outputs $y_{i}\in\mathbb{R}^{m}$, where $i\in\{1,\ldots,k\}$. Recall that we are training on $k$ samples with input dimension $n$ and output dimension $m$.
We denote the concatenated network outputs with network parameters $\theta\in\mathbb{R}^{d}$ using $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{km}$, where
We denote the concatenated target outputs as $y=\left[\;{y_{1}}^{\top},\;{y_{2}}^{\top},\;\ldots,\;{y_{k}}^{\top}\;\right]^{\top}\in\mathbb{R}^{km}$.
For any vector of target outputs $y\in\mathbb{R}^{km}$, we denote the set of parameters that fit them as
$\mathcal{Z}_{y}=\big\{\;\theta\in\mathbb{R}^{d}\;\big|\;\mathcal{F}(\theta)=y\;\big\}.$ (23)
Note that $\mathcal{Z}_{y}$ is exactly the zero-loss set with target outputs $y\in\mathbb{R}^{km}$.
We will show that $\mathcal{Z}_{y}$ contains no singular points for almost every $y\in\mathbb{R}^{km}$. More precisely, we show that $\mathcal{Z}_{y}\cap\mathrm{Crit}(\mathcal{F})=\emptyset$ for all $y\in\mathbb{R}^{km}$ except a set of Lebesgue measure zero.
We denote the set of target output vectors that lead to singularities as
$\xi\;=\;\big\{\;y\in\mathbb{R}^{km}\;\big|\;\mathcal{Z}_{y}\cap\mathrm{Crit}(\mathcal{F})\neq\emptyset\;\big\}$ (24)
Proposition B.1.
The set of target outputs that lead to singularities is exactly the image of the set of singular points:
$\xi\;=\;\mathcal{F}\big(\mathrm{Crit}(\mathcal{F})\big).$ (25)
Proof.
By our definitions, we have that $y\in\xi$ if and only if there exists $\theta\in\mathbb{R}^{d}$ such that $\mathcal{F}(\theta)=y$ and $\theta\in\mathrm{Crit}(\mathcal{F})$. First, any $\theta\in\mathrm{Crit}(\mathcal{F})$ implies that $\mathcal{F}(\theta)\in\xi$. Second, for some $y\in\mathbb{R}^{km}$, if no $\theta\in\mathrm{Crit}(\mathcal{F})$ exists such that $\mathcal{F}(\theta)=y$, then $y\not\in\xi$. ∎
Theorem B.2 (Morse–Sard theorem).
If a function $g:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ is continuously differentiable $k$ times, where $k\geq\mathrm{max}(a-b+1,1)$, then the image of its critical set $g\big(\mathrm{Crit}(g)\big)$ has Lebesgue measure zero in $\mathbb{R}^{b}$.
This result was first proved by Morse (1939) for the single-output functions (i.e., $b=1$), and later generalized to differentiable maps by Sard (1942).
By Assumption 4.2, the function $\mathcal{F}$ fits the continuous differentiability criteria. Therefore, we get that the set of target vectors that lead to singularities $\xi=\mathcal{F}\big(\mathrm{Crit}(\mathcal{F})\big)\subset\mathbb{R}^{km}$ has Lebesgue measure zero in $\mathbb{R}^{km}$.
Appendix C Proof of Theorem 4.14 (Gradient Orthogonality)
@@
Lemma C.1.
At a non-singular point of the zero-loss set, the tangent space is exactly the null space of the Hessian matrix, i.e. $T_{\theta}=\{\;v\in\mathbb{R}^{d}\;|\;\nabla^{2}\mathcal{L}(\theta)\,v=0\;\}\$for any non-singular $\theta\in\mathcal{Z}$.
@@
Part I: $\nabla^{2}\mathcal{L}(\theta)\;v=0\;\Rightarrow\,v\in T_{\theta}$.
@@
Theorem C.2 (Inverse Function Theorem).
Assume that $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{km}$ is a continuously differentiable function with $d\geq km$ and $\mathrm{rank}(\nabla\mathcal{F}(\theta))=km$ for some $\theta\in\mathbb{R}^{d}$. Then, for all $v\in R^{d}$ such that $\nabla\mathcal{F}(\theta)\,v=0$, there exists a smooth trajectory $s:R\rightarrow R^{d}$ such that $s(0)=\theta,\ s^{\prime}(0)=v,\,$ and $\mathcal{F}(s(t))=\mathcal{F}(\theta)\,$ for all $t\in\mathbb{R}$.
@@
Part II: $v\in T_{\theta}\;\Rightarrow\,\nabla^{2}\mathcal{L}(\theta)\;v=0$.
@@
Definition C.3 (Normal Space).
Let $\mathcal{N}_{\phi}=\{\,\alpha\in\mathbb{R}^{d}\;|\;\alpha^{\top}v=0,\;\forall v\in T_{\phi}\,\}$ denote the normal space at a point $\phi\in\mathcal{Z}$.
Proposition C.4.
@@
We give a proof by contradiction. Assume that there exists $v\in T_{\mathrm{proj}_{\mathcal{Z}}(\theta)}$ such that $v^{\top}\left(\theta-\mathrm{proj}_{\mathcal{Z}}(\theta)\right)\neq 0$. Then, by Definition 4.11, there must exist a smooth trajectory $f:\mathbb{R}\rightarrow\mathbb{R}^{d}$ with $f(0)=\mathrm{proj}_{\mathcal{Z}}(\theta),\ f^{\prime}(0)=v,$ and $f(t)\in\mathcal{Z}$ for all $t\in R$. By moving $\mathrm{proj}_{\mathcal{Z}}(\theta)$ along this trajectory, we can get closer to $\theta$. However, this should not be possible according to Definition 4.13. ∎
We now establish our main result: See 4.14
@@
with $\phi\in\mathcal{Z},\,\alpha\in\mathcal{N}_{\phi},\,x\in R^{+},$ and $v\in T_{\phi}$.
@@
for all $\phi\in S\cap\mathcal{Z},\,\alpha\in U(\mathcal{N}_{\phi}),\,x>0,\,v\in T_{\phi}\$, where $U(\mathcal{N}_{\phi})=\{\,v\in\,\mathcal{N}_{\phi}\;|\;\|v\|=1\,\}$.
@@
Since $v\in T_{\phi}$, from Lemma C.1, we have that $v^{\top}\,\nabla^{2}\mathcal{L}(\phi)\,\alpha=0$. This gives
@@
for any $x<a_{\inf}$. An appropriate choice of $C$ gives the desired bound for all $\theta$ with $\mathrm{dist}_{\mathcal{Z}}\theta<x_{0}$ for some $x_{0}>0$, for example: